# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@
Abstract
We construct a Lagrangian multiform for the class of cyclotomic (rational) Gaudin models by formulating its hierarchy within the Lie dialgebra framework of Semenov-Tian-Shansky and by using the framework of Lagrangian multiforms on coadjoint orbits. This provides the first example of a Lagrangian multiform for an integrable hierarchy whose classical $r$-matrix is non-skew-symmetric and spectral parameter-dependent. As an important by-product of the construction, we obtain a Lagrangian multiform for the periodic Toda chain by choosing an appropriate realisation of the cyclotomic Gaudin Lax matrix. This fills a gap in the landscape of Toda models as only the open and infinite chains had been previously cast into the Lagrangian multiform framework. A slightly different choice of realisation produces the so-called DST model. We demonstrate the versatility of the framework by coupling the periodic Toda chain with the DST model and by obtaining a Lagrangian multiform for the corresponding integrable hierarchy.
@@ 1 Introduction
A characteristic property of an integrable system is that its equations of motion can be seen as members of a hierarchy of compatible equations. Traditional Lagrangians fail to capture this notion of commuting (Hamiltonian) flows, or its discrete analog known as multidimensional consistency [N1, BS]. This obstacle to describing integrable hierarchies variationally was overcome for the first time in [LN] by introducing a new object — a Lagrangian multiform — together with a generalised variational principle applied to an appropriate generalisation of a classical action, and noticing a fundamental property of the multiform: the so-called closure relation. Since its introduction, this idea has been developed in several directions. Its connections with more traditional features of integrability (Lax pairs and Hamiltonian structures, for instance) have been established in the various realms of integrable systems: discrete and continuous finite-dimensional systems [YKLN, Su, PS, CDS], continuous infinite-dimensional systems – field theories in $1+1$ dimensions [SuV, SNC1, SNC2, CS1, CS2, PV, CSV] and in $2+1$ dimensions [SNC3, N2] – and semi-discrete systems [SV]. The relations between discrete and continuous multiforms were explored in [V]. The concept has even been extended to non-commuting flows in [CNSV].
In general, hierarchies of models in $D+1=d$ dimensions are described by a Lagrangian multiform which is a $d$-form integrated over a hypersurface of dimension $d$ in a so-called multi-time space of dimension greater than $d$ to yield an action functional depending not only on the field configurations but also on the hypersurface itself. This last point is the crucial new ingredient of the generalised variational principle used in Lagrangian multiform theory. One postulates a principle of least action which must be valid for any hypersurface embedded in the multi-time space. This captures the commutativity of the flows variationally and was adopted as a definition of pluri-Lagrangian systems, see [PS, PV] and references therein. In Lagrangian multiform theory, there is an additional postulate which leads to the closure relation. The latter has been shown [Su, CDS] to be equivalent to the Poisson involutivity of Hamiltonians, the Liouville criterion for integrability.
The generalised variational principle produces equations that come in two flavours: 1) Euler-Lagrange equations associated with each of the coefficients of the Lagrangian multiform which form a collection of Lagrangian densities; 2) Corner or structure equations on the Lagrangian coefficients themselves which select possible models and ensure the compatibility of the various equations of motion imposed on a common set of fields. This paper deals with a finite-dimensional system, so let us illustrate the ingredients for this case. The basic objects are a Lagrangian 1-form
$\mathscr{L}[q]=\sum_{k=1}^{N}\,\mathscr{L}_{k}[q]\,\mathrm{d}t_{k}$ (1.1)
@@
$S[q,\Gamma]=\int_{\Gamma}\mathscr{L}[q]$ (1.2)
where $\Gamma$ is a curve in the multi-time $\mathbb{R}^{N}$ with (time) coordinates $t_{1},\dots,t_{N}$ and $q$ denotes generic configuration coordinates. For instance, $q$ could be a position vector in $\mathbb{R}^{d}$ for some $d$, or as will be the case for us, an element of a (matrix) Lie group. The notations $\mathscr{L}[q]$ and $\mathscr{L}_{k}[q]$ mean that these quantities depend on $q$ and a finite number of derivatives of $q$ with respect to the times $t_{1},\dots,t_{N}$. In this paper, we restrict ourselves only to the case of first derivatives and simply write $\mathscr{L}_{k}$ for the Lagrangian coefficients. The application of the generalised variational principle leads to the following multi-time Euler-Lagrange equations [Su]
@@
$\frac{\partial\mathscr{L}_{k}}{\partial q_{t_{\ell}}}=0\,,\quad\ell\neq k\,,$ (1.4)
$\frac{\partial\mathscr{L}_{k}}{\partial q_{t_{k}}}=\frac{\partial\mathscr{L}_{\ell}}{\partial q_{t_{\ell}}}\,,\quad k,\ell=1,\dots,N\,.$ (1.5)
Note that (1.3) is simply the standard Euler-Lagrange equation for each $\mathscr{L}_{k}$. Condition (1.4) states that the Lagrangian coefficient $\mathscr{L}_{k}$ cannot depend on the velocities $q_{t_{\ell}}$ for $\ell\neq k$. The last condition (1.5) requires that the conjugate momentum to $q$ be the same with respect to all times $t_{k}$. The closure relation then stipulates that
$\mathrm{d}\mathscr{L}[q]=0\leavevmode\nobreak\ \leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \partial_{t_{k}}\mathscr{L}_{j}-\partial_{t_{j}}\mathscr{L}_{k}=0\,,$ (1.6)
on solutions of (1.3)-(1.5).
Gaudin models are a general class of integrable systems associated with Lie algebras with a nondegenerate invariant bilinear form. They were first introduced for the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ by M. Gaudin in [G1] as quantum integrable spin chains with long-range interactions, and then generalised for arbitrary semi-simple Lie algebras in [G2]. The integrable structure of this model, at both the classical and quantum levels, is underpinned by the rational skew-symmetric solution
@@
$[r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),r_{{}_{\mathbf{\underline{13}}}}(\lambda,\nu)]+[r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),r_{{}_{\mathbf{\underline{23}}}}(\mu,\nu)]+[r_{{}_{\mathbf{\underline{32}}}}(\nu,\mu),r_{{}_{\mathbf{\underline{13}}}}(\lambda,\nu)]=0\,,$ (1.8)
with $C_{{}_{\mathbf{\underline{12}}}}$ being the tensor Casimir of the underlying Lie algebra.
The cyclotomic Gaudin model of interest in this paper arises as a specialisation of a general procedure which can be traced back to reduction group ideas [M], first applied in the form of an averaging procedure [RF] to the rational classical $r$-matrix (1.7) in order to produce the trigonometric and elliptic $r$-matrices. This was generalised in various ways, for instance, in [AT, Av1, Av2] in the context of Sklyanin’s (linear) Poisson algebra
$\{U_{{}_{\mathbf{\underline{1}}}}(\lambda)\,,\,U_{{}_{\mathbf{\underline{2}}}}(\mu)\}=\left[r_{{}_{\mathbf{\underline{12}}}}^{0}(\lambda,\mu),\,U_{{}_{\mathbf{\underline{1}}}}(\lambda)+U_{{}_{\mathbf{\underline{2}}}}(\mu)\right].$ (1.9)
@@
$\phi:U(\lambda)\mapsto\omega\sigma^{-1}(U(\omega\lambda))$ (1.10)
where $\omega$ is a $T$th root of unity. Thus, $\phi$ is an automorphism of the Poisson algebra (1.9). This leads us to consider the fixed point subalgebra generated by
$L(\lambda)=\frac{1}{T}\sum_{k=0}^{T-1}\omega^{-k}\sigma^{k}U(\omega^{-k}\lambda)\,,\quad\sigma(L(\lambda))=\omega L(\omega\lambda)\,.$ (1.11)
@@
$\{L_{{}_{\mathbf{\underline{1}}}}(\lambda)\,,\,L_{{}_{\mathbf{\underline{2}}}}(\mu)\}=\left[r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)\,,\,L_{{}_{\mathbf{\underline{1}}}}(\lambda)\right]-\left[r_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda)\,,\,L_{{}_{\mathbf{\underline{2}}}}(\mu)\right].$ (1.12)
@@
is a non-skew-symmetric solution of CYBE. These facts are special cases of Propositions 2.2 and 4.1 in [CC]. It is well-known [BV] that the Poisson algebra (1.12) ensures that the quantities ${\rm Tr}L(\lambda)^{p}$ Poisson commute and the equations of motion generated by them with respect to (1.12) take the Lax form. More precisely, the generating Hamiltonian
$H(\mu)=\frac{1}{p+1}\operatorname{Tr}L(\mu)^{p+1}\,,\quad p\in\mathbb{Z}_{\geq 0}\,,$ (1.14)
satisfies $\{L(\lambda)\,,\,H(\mu)\}=\left[M(\lambda,\mu)\,,\,L(\lambda)\right]$ where
$M_{{}_{\mathbf{\underline{1}}}}(\lambda,\mu)=\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)L_{{}_{\mathbf{\underline{2}}}}(\mu)^{p}\,.$ (1.15)
@@
If one applies this construction to a Lax matrix of rational Gaudin model type, with poles in the finite set $D=\{0,\zeta_{1},\ldots,\zeta_{N},\infty\}\subset\mathbb{C}P^{1}$,
$U(\lambda)=\sum_{n=0}^{N_{0}-1}\frac{U_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{n=0}^{N_{r}-1}\frac{U_{r}^{(n)}}{(\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{N_{\infty}}U_{\infty}^{(n)}\lambda^{n}\,,\quad U_{s}^{(n)}\in\mathfrak{g}\,,$ (1.16)
then one obtains the so-called cyclotomic Gaudin model [S, VY1, VY2] and its associated integrable hierarchy. Note that in (1.16), we have anticipated that the pole at 00 and at infinity behave differently in (1.11) compared to the poles at $\lambda=\zeta_{r}\neq 0,\infty$. Although we will present the general algebraic setup, for simplicity, in our examples we will only consider the case where $U(\lambda)$ has simple poles at all $\zeta_{r}$, $r\in\{1,\ldots,N\}$, and double poles at the origin and at infinity. This will be sufficient for our application to the periodic Toda chain and the DST model. The corresponding Lax matrix of the cyclotomic Gaudin model then takes the form
@@
Our goal is to give a Lagrangian multiform description of the cyclotomic Gaudin hierarchy. In a nutshell, what is required is to interpret $L(\lambda)$ as an element of an appropriate coadjoint orbit so that we can write our Lagrangian multiform following the construction introduced in [CDS]. The latter is based on using the Lie dialgebra formalism [STS] which is reviewed in Section 2. In Section 3, we show how to construct the required algebraic ingredients for the cyclotomic Gaudin model. This is applied in Section 3.4 where we obtain a Lagrangian multiform for the cyclotomic Gaudin hierarchy and discuss its properties.
Section 4 deals with two special realisations of the cyclotomic Gaudin model: the periodic Toda chain (Section 4.1) and the integrable case of the discrete self–trapping (DST) model (Section 4.2). We derive Lagrangian multiforms for the hierarchies corresponding to these two theories using our results from Section 3. The construction of the periodic Toda multiform, in particular, complements the results of [CDS] where a Lagrangian multiform for the open Toda chain was derived, while in [SV] the infinite Toda chain was considered within a semi-discrete Lagrangian multiform. Finally, in Section 4.3, we illustrate how to couple the periodic Toda and DST hierarchies together in a straightforward manner and derive a Lagrangian multiform for this coupled hierarchy. We end with concluding remarks in Section 5.

### v2

# Lagrangian Multiform for Cyclotomic Gaudin Models

@@
[…]
Lagrangian Multiform for Cyclotomic Gaudin Models
[…]
Lagrangian Multiform for Cyclotomic Gaudin Models
[…]
Vincent CAUDRELIER a, Anup Anand SINGH a and Benoît VICEDO b
[…]
V. Caudrelier, A.A. Singh and B. Vicedo
[…]
Received May 28, 2024, in final form November 07, 2024; Published online November 15, 2024
[…]
We construct a Lagrangian multiform for the class of cyclotomic (rational) Gaudin models by formulating its hierarchy within the Lie dialgebra framework of Semenov-Tian-Shansky and by using the framework of Lagrangian multiforms on coadjoint orbits. This provides the first example of a Lagrangian multiform for an integrable hierarchy whose classical $r$-matrix is non-skew-symmetric and spectral parameter-dependent. As an important by-product of the construction, we obtain a Lagrangian multiform for the periodic Toda chain by choosing an appropriate realisation of the cyclotomic Gaudin Lax matrix. This fills a gap in the landscape of Toda models as only the open and infinite chains had been previously cast into the Lagrangian multiform framework. A slightly different choice of realisation produces the so-called discrete self-trapping (DST) model. We demonstrate the versatility of the framework by coupling the periodic Toda chain with the DST model and by obtaining a Lagrangian multiform for the corresponding integrable hierarchy.
[…]
Lagrangian multiforms; integrable systems; classical $r$-matrix; Gaudin models
[…]
17B80; 37J35; 70H06
@@ 1 Introduction
A characteristic property of an integrable system is that its equations of motion can be seen as members of a hierarchy of compatible equations. Traditional Lagrangians fail to capture this notion of commuting (Hamiltonian) flows, or its discrete analog known as multidimensional consistency [9, 25]. This obstacle to describing integrable hierarchies variationally was overcome for the first time in [23] by introducing a new object – a Lagrangian multiform – together with a generalised variational principle applied to an appropriate generalisation of a classical action, and noticing a fundamental property of the multiform: the so-called closure relation. Since its introduction, this idea has been developed in several directions. Its connections with more traditional features of integrability (Lax pairs and Hamiltonian structures, for instance) have been established in the various realms of integrable systems: discrete and continuous finite-dimensional systems [11, 27, 36, 43], continuous infinite-dimensional systems – field theories in $1+1$ dimensions [13, 14, 15, 28, 32, 33, 37] and in $2+1$ dimensions [26, 34] – and semi-discrete systems [35]. The relations between discrete and continuous multiforms were explored in [40]. The concept has even been extended to non-commuting flows in [12].
In general, hierarchies of models in $d$ spacetime dimensions are described by a Lagrangian multiform which is a $d$-form integrated over a hypersurface of dimension $d$ in a so-called multi-time space of dimension greater than $d$ to yield an action functional depending not only on the field configurations but also on the hypersurface itself. This last point is the crucial new ingredient of the generalised variational principle used in Lagrangian multiform theory. One postulates a principle of least action which must be valid for any hypersurface embedded in the multi-time space. This captures the commutativity of the flows variationally and was adopted as a definition of pluri-Lagrangian systems, see [27, 28] and references therein. In Lagrangian multiform theory, there is an additional postulate which leads to the closure relation. The latter has been shown [11, 36] to be equivalent to the Poisson involutivity of Hamiltonians, the Liouville criterion for integrability.
The generalised variational principle produces equations that come in two flavours: 1) Euler–Lagrange equations associated with each of the coefficients of the Lagrangian multiform which form a collection of Lagrangian densities; 2) Corner or structure equations on the Lagrangian coefficients themselves which select possible models and ensure the compatibility of the various equations of motion imposed on a common set of fields. This paper deals with a finite-dimensional system, so let us illustrate the ingredients for this case. The basic objects are a Lagrangian 1-form
$\mathscr{L}[q]=\sum_{k=1}^{N}\mathscr{L}_{k}[q]\,\mathrm{d}t_{k}$
@@
$S[q,\Gamma]=\int_{\Gamma}\mathscr{L}[q],$
where $\Gamma$ is a curve in the multi-time $\mathbb{R}^{N}$ with (time) coordinates $t_{1},\dots,t_{N}$, and $q$ denotes generic configuration coordinates. For instance, $q$ could be a position vector in $\mathbb{R}^{d}$ for some $d$, or as will be the case for us, an element of a (matrix) Lie group. The notations $\mathscr{L}[q]$ and $\mathscr{L}_{k}[q]$ mean that these quantities depend on $q$ and a finite number of derivatives of $q$ with respect to the times $t_{1},\dots,t_{N}$. In this paper, we restrict ourselves only to the case of first derivatives and simply write $\mathscr{L}_{k}$ for the Lagrangian coefficients. The application of the generalised variational principle leads to the following multi-time Euler–Lagrange equations [36]:
@@
$\frac{\partial\mathscr{L}_{k}}{\partial q_{t_{\ell}}}=0,\qquad\ell\neq k,$ (1.2)
$\frac{\partial\mathscr{L}_{k}}{\partial q_{t_{k}}}=\frac{\partial\mathscr{L}_{\ell}}{\partial q_{t_{\ell}}},\qquad k,\ell=1,\dots,N.$ (1.3)
Note that (1.1) is simply the standard Euler–Lagrange equation for each $\mathscr{L}_{k}$. Equations (1.2) and (1.3) are the corner equations. It turns out that these corner equations will, in fact, be identically satisfied by our Lagrangian multiforms, but in general they can represent non-trivial equations for $q$. The closure relation then stipulates that
$\mathrm{d}\mathscr{L}[q]=0\quad\Leftrightarrow\quad\partial_{t_{k}}\mathscr{L}_{j}-\partial_{t_{j}}\mathscr{L}_{k}=0$
on solutions of (1.1)–(1.3).
Gaudin models are a general class of integrable systems associated with Lie algebras with a nondegenerate invariant bilinear form. They were first introduced for the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ by M. Gaudin in [19] as quantum integrable spin chains with long-range interactions, and then generalised for arbitrary semi-simple Lie algebras in [20]. At both the classical and quantum levels, the integrable structure of this model, associated with a Lie algebra $\mathfrak{g}$, is underpinned by the rational skew-symmetric solution
@@
$\bigl{[}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),r_{{}_{\mathbf{\underline{13}}}}(\lambda,\nu)\bigr{]}+\bigl{[}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),r_{{}_{\mathbf{\underline{23}}}}(\mu,\nu)\bigr{]}+\bigl{[}r_{{}_{\mathbf{\underline{32}}}}(\nu,\mu),r_{{}_{\mathbf{\underline{13}}}}(\lambda,\nu)\bigr{]}=0.$
The rational classical $r$-matrix (1.4) takes values in $\mathfrak{g}\otimes\mathfrak{g}$ and depends on spectral parameters $\lambda,\mu\in\mathbb{C}$, and the tensor Casimir $C_{{}_{\mathbf{\underline{12}}}}\in\mathfrak{g}\otimes\mathfrak{g}$ appearing in (1.4) is defined as
$C_{{}_{\mathbf{\underline{12}}}}=I_{a}\otimes I^{a},$
where $\{I_{a}\}$ and $\{I^{a}\}$ are dual bases of $\mathfrak{g}$ with respect to a fixed nondegenerate invariant bilinear form. The main property of $C_{{}_{\mathbf{\underline{12}}}}$ which gives it its name is $\bigl{[}C_{{}_{\mathbf{\underline{12}}}},X_{{}_{\mathbf{\underline{1}}}}+X_{{}_{\mathbf{\underline{2}}}}\bigr{]}=0$ for all $X\in\mathfrak{g}$ and where we use the standard tensorial notation $X_{{}_{\mathbf{\underline{1}}}}\equiv X\otimes\mathbb{1}$ and $X_{{}_{\mathbf{\underline{2}}}}\equiv\mathbb{1}\otimes X$.
The cyclotomic Gaudin model of interest in this paper arises as a specialisation of a general procedure which can be traced back to reduction group ideas [24], first applied in the form of an averaging procedure [29] to the rational classical $r$-matrix (1.4) in order to produce the trigonometric and elliptic $r$-matrices. This was generalised in various ways, for instance, in [4, 5, 6] in the context of Sklyanin’s (linear) Poisson algebra
$\{U_{{}_{\mathbf{\underline{1}}}}(\lambda),U_{{}_{\mathbf{\underline{2}}}}(\mu)\}=\bigl{[}r_{{}_{\mathbf{\underline{12}}}}^{0}(\lambda,\mu),U_{{}_{\mathbf{\underline{1}}}}(\lambda)+U_{{}_{\mathbf{\underline{2}}}}(\mu)\bigr{]},$ (1.5)
where, as above, the index denotes which factor in the tensor product $\mathfrak{g}\otimes\mathfrak{g}$ the $\mathfrak{g}$-valued function $U$ sits in, that is, $U_{{}_{\mathbf{\underline{1}}}}(\lambda)\equiv U(\lambda)\otimes\mathbb{1}$ and $U_{{}_{\mathbf{\underline{2}}}}(\mu)\equiv\mathbb{1}\otimes U(\mu)$.
@@
$\phi\colon\ U(\lambda)\mapsto\omega\sigma^{-1}(U(\omega\lambda)),$
where $\omega$ is a $T$-th root of unity. Thus, $\phi$ is an automorphism of the Poisson algebra (1.5). This leads us to consider the fixed point subalgebra generated by
$L(\lambda)=\frac{1}{T}\sum_{k=0}^{T-1}\omega^{-k}\sigma^{k}U\bigl{(}\omega^{-k}\lambda\bigr{)},\qquad\sigma(L(\lambda))=\omega L(\omega\lambda).$ (1.6)
@@
$\bigl{\{}L_{{}_{\mathbf{\underline{1}}}}(\lambda),L_{{}_{\mathbf{\underline{2}}}}(\mu)\bigr{\}}=\bigl{[}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),L_{{}_{\mathbf{\underline{1}}}}(\lambda)\bigr{]}-\bigl{[}r_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),L_{{}_{\mathbf{\underline{2}}}}(\mu)\bigr{]},$ (1.7)
@@
is a non-skew-symmetric solution of CYBE. These facts are special cases of [10, Propositions 2.2 and 4.1]. It is well known [8] that the Poisson algebra (1.7) ensures that the quantities ${\rm Tr}L(\lambda)^{p}$ Poisson commute and the equations of motion generated by them with respect to (1.7) take the Lax form. More precisely, the generating Hamiltonian
$H(\mu)=\frac{1}{p+1}\operatorname{Tr}L(\mu)^{p+1},\qquad p\in\mathbb{Z}_{\geq 0},$
satisfies
$\{L(\lambda),H(\mu)\}=[M(\lambda,\mu),L(\lambda)],$
where by definition $\{L(\lambda),H(\mu)\}=\{L^{a}(\lambda),H(\mu)\}I_{a}$, and
$M_{{}_{\mathbf{\underline{1}}}}(\lambda,\mu)=\operatorname{Tr}_{{}_{\mathbf{\underline{2}}}}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)L_{{}_{\mathbf{\underline{2}}}}(\mu)^{p}.$
@@
If one applies this construction to a Lax matrix of rational Gaudin model type, with poles in the finite set $D=\{0,\zeta_{1},\dots,\zeta_{N},\infty\}\subset\mathbb{C}P^{1}$,
$U(\lambda)=\sum_{n=0}^{N_{0}-1}\frac{U_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{n=0}^{N_{r}-1}\frac{U_{r}^{(n)}}{(\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{N_{\infty}}U_{\infty}^{(n)}\lambda^{n},\qquad U_{s}^{(n)}\in\mathfrak{g},$ (1.9)
then one obtains the so-called cyclotomic Gaudin model [31, 41, 42] and its associated integrable hierarchy. Note that in (1.9), we have anticipated that the pole at 00 and at infinity behave differently in (1.6) compared to the poles at $\lambda=\zeta_{r}\neq 0,\infty$. Although we will present the general algebraic setup, for simplicity, in our examples we will only consider the case where $U(\lambda)$ has simple poles at all $\zeta_{r}$, $r\in\{1,\dots,N\}$, and double poles at the origin and at infinity. This will be sufficient for our application to the periodic Toda chain and the discrete self-trapping (DST) model. The corresponding Lax matrix of the cyclotomic Gaudin model then takes the form
@@
Our goal is to give a Lagrangian multiform description of the cyclotomic Gaudin hierarchy. In a nutshell, what is required is to interpret $L(\lambda)$ as an element of an appropriate coadjoint orbit so that we can write our Lagrangian multiform following the construction introduced in [11]. The latter is based on using the Lie dialgebra formalism [30] which is reviewed in Section 2. In Section 3, we show how to construct the required algebraic ingredients for the cyclotomic Gaudin model. This is applied in Section 3.4, where we obtain a Lagrangian multiform for the cyclotomic Gaudin hierarchy and discuss its properties.
Section 4 deals with two special realisations of the cyclotomic Gaudin model: the periodic Toda chain (Section 4.1) and the integrable case of the DST model (Section 4.2). We derive Lagrangian multiforms for the hierarchies corresponding to these two theories using our results from Section 3. The construction of the periodic Toda multiform, in particular, complements the results of [11] where a Lagrangian multiform for the open Toda chain was derived, while in [35] the infinite Toda chain was considered within a semi-discrete Lagrangian multiform. Finally, in Section 4.3, we illustrate how to couple the periodic Toda and DST hierarchies together in a straightforward manner and derive a Lagrangian multiform for this coupled hierarchy. We end with concluding remarks in Section 5.